OpenStudy (anonymous):

How do I find the x intercept of y=ln(x+3)

OpenStudy (anonymous):

To find y intercepts, set x equal to 0

OpenStudy (anonymous):

Its x intercept i know you do 0=ln(x+3) but how do i solve for x

OpenStudy (asnaseer):

use the log rule:\[\log_ab=c\implies b=a^c\]

OpenStudy (anonymous):

x+3=e^0?

OpenStudy (asnaseer):

perfect!

OpenStudy (anonymous):

Wow thats surprisingly simple. Thank you

OpenStudy (asnaseer):

yw :)